OpenStudy (anonymous):

can someone show me how to do this if m angle dab=2x-9 and m angle bdc=7x-27 solve for m angle dab

OpenStudy (anonymous):

can someone please help explain how to do this

jhonyy9 (jhonyy9):

suppose abcd wann being a rectangle ,yes so then you know that a rectangle has angles equal 90 degree so this mean that 2x-9=90 and 7x-27=90 2x-9=90 2x=99 x=99/2 7x-27=90 7x=117 x=117/2 or you can writing dab = bdc 2x-9=7x-27 -9+27=7x-2x 18=5x x=18/5 so do you can rewrite the exactle text of exercise please because i think you have missed anything